116grams of KF is dissolved in enough water to make 4.0 L of solution.
1
Expert's answer
2020-11-12T06:31:35-0500

m (KF) = 116 g

n (KF) = m (KF)/M (KF) = 116 / (39 + 19) = 58 / 116 = 0.5 mol

C (KF) = n (KF) / V(solution) = 0.5 mol / 4 L = 0.125 mol/L